Use the compound interest formula A=P (1+r/n)^nt to answer the following questions. In the formula, A is the amount of money in the savings account, P is the principal, r is the interest rate, t is the time that the money is in the account, and n is the number of times the money is compounded per year. Write each answer to the following questions in a complete sentence.
a. What will be the amount in the account if $800 is invested at 2.4% compounded monthly for 5 years?

b. What amount should a person place in an account if she wishes to have $1500 in 4 years if the money is compounded quarterly at an interest rate of 3.5%?
Thank you for your help!!

Answer by stanbon(75887)

a. What will be the amount in the account if $800 is invested at 2.4% compounded monthly for 5 years?
A=P(1+r/n)^(nt)
A=800(1+0.024/12)^(12*5)
A=800(1.002)^60
A=800*1.12736
A=$901.89
----------------
b. What amount should a person place in an account if she wishes to have $1500 in 4 years if the money is compounded quarterly at an interest rate of 3.5%?
1500=P(1+0.035/4)^(4*4)
1500=P(1.00875)^16
1500=P(1.14957
P=$1304.83
Cheers,
Stan H.